Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Bacon goes for the prize

Expert replies
by sanju09 » Tue Feb 21, 2012 4:46 am
At a carnival, there is a prize on randomly drawing a green ball in three fresh attempts, from a bag that contains 4 red, 3 green, and 2 white all identical balls. Bacon goes for the prize. What is the probability that he’d get the prize?
A. 4/27
B. 2/9
C. 1/3
D. 5/9
E. 19/27



[spoiler]made up by Sanjeev K Saxena for Avenues Abroad[/spoiler]
The mind is everything. What you think you become. –Lord Buddha



Sanjeev K Saxena
Quantitative Instructor
The Princeton Review - Manya Abroad
Lucknow-226001

www.manyagroup.com
Join the discussion
Source: — Problem Solving |

by MBACenter » Tue Feb 21, 2012 5:19 am
sanju09 wrote:At a carnival, there is a prize on randomly drawing a green ball in three fresh attempts, from a bag that contains 4 red, 3 green, and 2 white all identical balls. Bacon goes for the prize. What is the probability that he’d get the prize?
A. 4/27
B. 2/9
C. 1/3
D. 5/9
E. 19/27



[spoiler]made up by Sanjeev K Saxena for Avenues Abroad[/spoiler]
Depends. Do you mean the probability that he gets three green balls, or the probability that with three tries he gets at least one green ball? And do the marbles stay out, or does he put them back after each try?

If it's all three and the marbles stay out, then that would be:

(3/9) * (2/8) * (1/7) = (1/3) * (1/4) * (1/7) = 1/84.

If it's at least one of the three and the marbles stay out, then that would be one minus the probability that he DOESN'T pick any greens:

1 - (6/9) * (5/8) * (4/7) = 1 - (1/3) * (5/1) * (1/7) = 1 - 5/21 = 16/21

If it's all three and the marbles do not stay out, then that would be:

(3/9) * (3/9) * (3/9) = (1/3) * (1/3) * (1/3) = 1/27.

If it's at least one of the three and the marbles do not stay out, then that would be one minus the probability that he DOESN'T pick any greens:

1 - (6/9) * (6/9) * (6/9) = 1 - (2/3) * (2/3) * (2/3) = 1 - 8/27
Academic Coordinator
MBA Center Paris
Join the discussion

by sanju09 » Wed Feb 22, 2012 12:27 am
MBACenter wrote:
sanju09 wrote:At a carnival, there is a prize on randomly drawing a green ball in one of the three fresh attempts, from a bag that contains 4 red, 3 green, and 2 white all identical balls. Bacon goes for the prize. What is the probability that he’d get the prize?
A. 4/27
B. 2/9
C. 1/3
D. 5/9
E. 19/27



[spoiler]made up by Sanjeev K Saxena for Avenues Abroad[/spoiler]
Depends. Do you mean the probability that he gets three green balls, or the probability that with three tries he gets at least one green ball? And do the marbles stay out, or does he put them back after each try?

If it's all three and the marbles stay out, then that would be:

(3/9) * (2/8) * (1/7) = (1/3) * (1/4) * (1/7) = 1/84.

If it's at least one of the three and the marbles stay out, then that would be one minus the probability that he DOESN'T pick any greens:

1 - (6/9) * (5/8) * (4/7) = 1 - (1/3) * (5/1) * (1/7) = 1 - 5/21 = 16/21

If it's all three and the marbles do not stay out, then that would be:

(3/9) * (3/9) * (3/9) = (1/3) * (1/3) * (1/3) = 1/27.

If it's at least one of the three and the marbles do not stay out, then that would be one minus the probability that he DOESN'T pick any greens:

1 - (6/9) * (6/9) * (6/9) = 1 - (2/3) * (2/3) * (2/3) = 1 - 8/27
I meant this only, please suggest the adequate wordings.
Thanks MBACenter for pointing out the flaw. The term "fresh attempts" should mean that the experiments are carried out with replacement, if I am not wrong.
The mind is everything. What you think you become. –Lord Buddha



Sanjeev K Saxena
Quantitative Instructor
The Princeton Review - Manya Abroad
Lucknow-226001

www.manyagroup.com
Join the discussion